A stone is dropped from rest from a building that is 100 meters tall.
a) Find the velocity after 2.64 s. Incorrect: Your answer is incorrect. m/s
(b) Find the speed of the stone after same time as in part (a). m/s
(c) Find the vertical displacement after the same time as in part (a). m
(d) Sketch the vy versus t graph out to the same time as in part (a). Calculate the area "under" the graph. m
(e) At what time does the stone hit the ground?
1
Expert's answer
2019-02-17T09:43:07-0500

a) Velocity is negative (downwards).

V=-v=-gt=-(9.8)(2.64)=-25.9 m/s.

(b)

v=gt=(9.8)(2.64)=25.9 m/s.

(c)

D=-(gt^2)/2=-(9.8(2.64)^2)/2=-34.2 m.

(d)

v_y=gt=-9.8t.

Area=-1/2 (gt)(t)=-(gt^2)/2=-(9.8(2.64)^2)/2=-34.2 m.